OpenStudy (anonymous):

Which set of measurements will give you a volume of 168 cubic meters?

OpenStudy (anonymous):

base = 3 cm, height of base = 4 cm, height of prism = 7 cm base = 3 cm, height of base = 8 cm, height of prism = 7 cm base = 6 cm, height of base = 4 cm, height of prism = 7 cm base = 6 cm, height of base = 8 cm, height of prism = 7 cm

OpenStudy (campbell_st):

which 3 numbers multiply to 168 I know that the divisibility test for 3 is add the digits... if the answer is a multiply of 3 then 3 will be a factor so 1 + 6 + 8 =15 so 3 is a factor 168/3 = 56 now find te factors of 56... and then your answer will be 3, and the factors of 56

OpenStudy (campbell_st):

if 3 is a factor and the number is even then 6 is a factor 168/6 = 28 find the factors of 28 so the answer could be, 6, and the factors of 28